The curve formed by a rope bridge can be modeled by the relation y = x^2 - 11x + 10, where "x" is the horizontal distance in metres and "y" is the height in metres. What is the horizontal distance from one end of the bridge to the other end?


show me how

1
Expert's answer
2021-12-15T16:57:01-0500
"y=0=>x^2 - 11x + 10=0"

"(x-1)(x-10)=0"

"x_1=1, x_2=10"

"distance=d=x_2-x_1=10-1=9(m)"

The horizontal distance from one end of the bridge to the other end is "9" metres.